The instantaneous power delivered to an object is P 4t2 W. If the object has a force of 3 s? 6 N acting in its direction of velocity, what is the instantaneous speed of the object at t A. 114 m/s B. 6 m/s C. 2 m/s D. 1/6 m/s E. 4 m/s MacBook Pro s lock command option control optioncommand
Answer #1

P=4t^2, F=6N, t=3s, v=?

P=Fv,\:\ v=P/F

v=\frac{4t^2}{6}=\frac{2t^2}{3}

At t=3s Instantaneous speed is

v=\frac{2\times 3^2}{3}=6m/s